OpenStudy (anonymous):

Use the given conditions to write an equation for the line in point-slope form and slope-intercept form. Passing through (6,91) and (−8,7).

OpenStudy (aravindg):

use two point form

OpenStudy (anonymous):

m= (x2-x1)/(y2-y1)?

OpenStudy (aravindg):

no take a look at this tutorial http://easycalculation.com/analytical/learn-two-point.php

OpenStudy (anonymous):

^this is exactly what I did. I ended up with the answer:y= (21/2)x+15 and it's wrong.

OpenStudy (anonymous):

i'm redoing it i think i see my problem.

OpenStudy (anonymous):

My answer this time is (21/2)x + 28 and I don't see where I'm going wrong.

OpenStudy (dumbcow):

your slope is wrong \[m = \frac{y_{2} -y_{1}}{x_{2} -x_{1}}\] change in 'y' is on top

OpenStudy (aravindg):

yep :)

OpenStudy (anonymous):

yeah i noticed that after i typed it in.but i have been doing your formula the whole time on my whiteboard and still getting it wrong

OpenStudy (dumbcow):

ok, using the given points \[m = \frac{84}{14} = 6\] then \[y - 7 = 6(x-(-8))\] \[y-7 = 6x+48\] \[y=6x+55\]

OpenStudy (raden_zaikaria):